OpenStudy (anonymous):

What is the area of the parallelogram? A parallelogram is shown. The bottom side is labeled sixteen inches. A dotted line extends from the lower right corner of the parallelogram to the right. This line intersects a second dotted line that stretches from the upper right corner of the parallelogram downward, intersecting the first dotted line to form a right angle. This dotted line is labeled fourteen inches. sixty square inches one hundred ninety-six square inches two hundred twenty-four square inches two hundred fifty-six square inches

OpenStudy (anonymous):

@crewfan and @arabpride please help me with this

OpenStudy (anonymous):

fan and medal of you already aren't fanded by me LOL i can't spell

OpenStudy (anonymous):

Please help me

OpenStudy (anonymous):

how tall is the parallelogram?

OpenStudy (arabpride):

Would u mind drawing it out? *^.^*

OpenStudy (anonymous):

 Parallelograms are kinda like rectangles, in the sense that both their areas satisfies the equation A=b*h. I believe my interpretation is correct. The vertical dotted line 14 in, and the solid bottom line is 16 in. But the vertical dotted line is the height, and the solid line is the base! The area is the same as a rectangle of those dimensions, because if you can imagine cutting a piece of the parallelogram on the left, like so:  The pieces fit perfectly into a rectangle! So the answer is A = b*h = 16*14 = 224!